Given f(x)=3x+1 and g(x)=1/2, what is the domain of f,g & fog
Answer by venugopalramana(3286)

Given f(x)=3x+1 and g(x)=1/2, what is the domain of f,g & fog
DOMAIN OF F(X) IS ALL REAL NUMBERS
DOMAIN OF G(X) IS ALL REAL NUMBERS
FOG=3*1/2+1=2.5
ITS DOMAIN IS ALL REAL NUMBERS